# Advances in Feed-Forward 3D Reconstruction and View Synthesis: A Survey

### v2

@@ Abstract
Figure 1: A summary of representative feed-forward works based on their category and timeline.
Figure 2: This survey discusses the methodology and application of feed-forward models for various 3D reconstruction and Novel View Synthesis (NVS) tasks as listed in the figure. The samples are adapted from [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
@@ 1 Introduction
3D reconstruction and rendering are long-standing and central challenges in computer vision and computer graphics. They enable a wide range of applications, from digital content creation, augmented reality, and virtual reality to robotics, autonomous systems, and digital twins. Traditionally, high-quality 3D reconstruction and view synthesis has relied on optimization-based pipelines such as Structure-from-Motion (SfM) and Multi-View Stereo (MVS). However, these methods are often computationally expensive, slow to converge, and dependent on precisely calibrated datasets, limiting their practicality in real world scenarios. In light of this, feed-forward methods emerged as an important research line in 3D vision.
@@
Feef-forward models have been studied for a long time, such as early works on cost-volume-based Multi-View Stereo [13] and layered representations such as multiplane images (MPI) [14]. These methods demonstrated the potential of learning-based inference without per-scene optimization. In recent years, fueled by breakthroughs in deep learning and neural representations, feed-forward methods [15, 16] have emerged as a transformative alternative in 3D reconstruction and view synthesis. Unlike classical methods that require iterative optimization per scene, feed-forward models infer 3D geometry or novel views in a single forward pass, enabling orders of magnitude faster inference with improved generalization. These models leverage learned priors and large-scale training to make predictions, making them especially appealing for time-sensitive and scalable applications, such as robotic perception and interactive 3D asset creation.
@@
This survey focuses on feed-forward methods developed primarily after the emergence of neural radiance fields (NeRF) [17] in 2020, which catalyzed a rapid evolution in feed-forward fashion as shown in Fig. 1. We present a comprehensive review of feed-forward methods for 3D reconstruction and view synthesis, with an emphasis on the core architectures, scene representations, and downstream applications that define this fast-evolving area. We systematically categorize existing approaches based on their underlying scene representations, which determine how 3D structure and appearance are modeled and rendered. Specifically, we identify five major categories: 1) models built on Neural Radiance Fields (NeRF) [17], which leverage volumetric rendering through learned radiance fields; 2) pointmap-based approaches [1], which operate on pixel-aligned 3D pointmaps; 3) 3D Gaussian Splatting (3DGS)-based models [18], which use rasterizable Gaussian primitives for fast and efficient rendering; 4) methods based on other 3D representations like mesh, occupancy and signed distance function (SDF); and 5) 3D-representation-free models, which leverage deep neural networks to synthesize views directly without a explicit 3D representation. For each category, we provide an in-depth analysis of representative and state-of-the-art methods, highlighting their core architectural designs, feature representations, and the inductive biases embedded in their formulations.
@@
We also discuss high-impact 3D vision applications enabled by feed-forward methods as shown in Fig. 2, which provide scalable, fast, and generalizable solutions across domains. These include pose-free and dynamic 3D reconstruction, 3D-aware image and video synthesis, and camera-controllable video generation. Additionally, these models facilitate semantic reasoning and dense matching, advancing tasks such as 3D-aware segmentation and optical flow estimation. In robotics and SLAM, feed-forward models enable real-time scene understanding and tracking, while in digital humans, they support efficient yet generalizable avatar reconstruction from sparse inputs.
@@
Despite impressive progress, feed-forward models still face major challenges, including limited modality diversity in datasets, poor generalization in free-viewpoint synthesis, and the high computational cost of long-context processing. Addressing these challenges will require advances in efficient architectures and scalable datasets. Finally, we conclude with the societal impact of this technology, highlighting the importance of responsible deployment and transparent modeling practices.
@@ 2 Methods
In the following, we broadly categorize the feed-forward 3D reconstruction and view synthesis methods into five categories based on their underlying representation: NeRF models (Sec. 2.1), Pointmap models (Sec. 2.2), 3DGS models (Sec. 2.3), models employing other common representations (e.g., mesh, occupancy, SDFs in Sec. 2.4), and 3D representation-free models (Sec. 2.5).
Figure 3: Representative frameworks of feed-forward NeRF. The samples are adapted from  [19] [20] [21] [10].
@@ 2.1 NeRF
Neural radiance fields (NeRF) [17] have recently gained significant attention for high-quality novel view synthesis using implicit scene representations and differentiable volume rendering. By leveraging MLPs, NeRF reconstructs 3D scenes from multiview 2D images, enabling the generation of novel views with excellent multiview consistency. However, a major limitation of NeRF is its requirement for per-scene optimization, which restricts its generalization to unseen scenes. To address this, feed-forward approaches have been proposed, where neural networks learn to infer NeRF representations directly from sparse input views, thereby eliminating the need for scene-specific optimization. As a pioneering feed-forward NeRF work, PixelNeRF [22] introduces a conditional NeRF framework that leverages pixel-aligned image features extracted from input images, allowing the model to generalize across diverse scenes and perform novel view synthesis from sparse observations. A large number of follow-ups adopt various techniques for feed-forward NeRF, and we broadly categorize them into the following categories based on feature representations.
@@
Several methods have been proposed to encode a global 1D latent code for NeRF prediction, where the same latent code is shared between all 3D points in a scene. For example, CodeNeRF [19], as illustrated in Fig. 3 (a), introduces a disentanglement strategy that jointly learns separate embeddings for texture and shape, along with an MLP conditioned on these embeddings to predict the color and volumetric density of each 3D point. ShaRF [23] introduces latent codes for shape and appearance, which serve as conditioning inputs for NeRF reconstruction. In addition, Shap-E [24], following Point-E [25], encodes point clouds and RGBA input images into a series of latent vectors, which are subsequently utilized for NeRF prediction.
@@
2D feature-based methods typically leverage an image encoder to extract image features of source views and obtain features of arbitrary 3D points by ray projection without relying on 3D intermediate features. For example, GRF [20], as illustrated in Fig. 3(b) projects each 3D point along a camera ray onto source views to extract corresponding multiview features. These features are then aggregated and passed through an MLP to predict RGB color and volumetric density. IBRNet [26] follows a comparable approach, projecting 3D points onto nearby source views to extract image features that are aggregated across views for radiance field inference. NeRFormer [27] also employs ray-projected features and performs multiview feature aggregation to guide the NeRF prediction. Besides, SRF [28] projects 3D points onto multiview input images to construct a stereo feature matrix, which is processed by a 2D CNN to produce view-aligned features for color and density prediction. To provide additional geometric cues for color and density prediction, GNT [29] introduces a view transformer that leverages epipolar constraint to aggregate projected features from multiple views in a geometrically consistent manner. Its successor, GNT-MOVE [30], bridges the view transformer with the Mixture-of-Experts concept from large language models, enhancing its cross-scene generalization capability. MatchNeRF [31] explicitly models the correspondence information by computing the similarity between ray-projected features from pairs of nearby source views, using this information as a conditioning input for the prediction. ContraNeRF [32] introduces geometry-aware feature extraction and contrastive learning [33] to query features from multiple source views and aggregate them to obtain geometrically enhanced feature maps.
@@
3D Volume Features. MVSNeRF [21] is inspired by multi-view stereo (MVS) [13, 34, 35] and constructs cost volumes from input images as shown in Fig. 3(c). These cost volumes are used to generate a neural scene encoding volume that stores per-voxel features capturing both local geometry and appearance. For any 3D point, its features are obtained via trilinear interpolation from the encoding volume and then decoded by an MLP to predict the corresponding density and color. To improve rendering quality in both fine-detail areas and occluded regions, GeoNeRF [36] extends MVSNeRF by first constructing cascaded cost volumes for each source view, followed by an attention-based volume aggregation across views. NeuRay [37] is also proposed to address the issue of occlusion by leveraging constructed cost volumes to predict the visibility of 3D points, which can identify feature inconsistencies caused by occlusion. WaveNeRF [38] designs a wavelet multiview stereo that incorporates wavelet frequency volumes into the MVS to preserve high-frequency information and achieve desirable scene geometry reconstruction. In addition, for efficient rendering, ENeRF [39] proposes sampling a limited number of points near the scene surface by predicting the coarse scene geometry from a constructed cascade cost volume, enabling improved rendering speed. MuRF [40] eliminates the use of cost volumes for predefined reference input views, instead constructing a target view frustum volume to effectively aggregate information from the input images, particularly in scenes with limited overlap between the reference and target views. To improve the quality of geometry estimation, GeFu [41] introduces an adaptive cost aggregation module that reweights the contributions of different source views, allowing the model to learn adaptive weights for constructing cost volumes.
@@
3D Triplane Features. The triplane serves as an efficient volumetric representation [42, 43], making it highly compatible with feed-forward models. Specifically, Large Reconstruction Model (LRM) [10] employs a large transformer-based encoder-decoder architecture and directly regresses a feature triplane representation as shown in Fig. 3(d), enabling NeRF prediction from triplane features. Pf-LRM [44] extends the LRM into a pose-free setting, which jointly reconstructs the triplane NeRF representations and predicts relative camera poses. TripoSR [45] further enhances LRM through improvements in data curation and rendering, model architecture, and training strategies. Considering the scarcity, licensing constraints, and inherent biases of 3D data, LRM-Zero [46] is proposed to enable training solely on synthesized data from Zeroverse [46]. In addition, several methods combine a large reconstruction model with a diffusion model. For example, Instant3D [47] first leverages a fine-tuned 2D diffusion model [48] to generate 4-view images from a text prompt and then uses a transformer-based large reconstruction model to predict a NeRF. DMV3D [49], inspired by RenderDiffusion [50], incorporates LRM into multiview diffusion, which gradually reconstructs a clean triplane NeRF representation from noisy multiview images in the diffusion process.
@@
In addition to the aforementioned methods, several efforts have also focused on feed-forward NeRF reconstruction with other types of features. For example, VisionNeRF [51] proposes to leverage vision transformer [52] and convolutional networks to extract global 1D features and 2D image features, respectively, and constructs a multi-level feature map that serves as the conditioning inputs of NeRF prediction to enhance rendering quality, particularly in occluded regions. MINE [53] integrates NeRF and multiplane image (MPI) [54] representations to enable generalizable, occlusion-aware 3D reconstruction from a single image.
@@ 2.2 Pointmap
Figure 4: The framework of feed-forward pointmap reconstruction. The samples are adapted from [3].
@@
To handle more views, Fast3R [59] builds on DUSt3R and designs a global fusion transformer to process multiview inputs simultaneously. MV-DUSt3R [60] instead leverages multiview decoder blocks to learn both reference-to-source and source-to-source view relationships, thereby extending DUSt3R to a multiview setting. SLAM3R [61] introduces an Image-to-Points module that enables simultaneous processing of multiview inputs, effectively enhancing reconstruction quality without sequential reconstruction.
@@
To improve computational efficiency, a couple of works introduce a memory mechanism that aims to incrementally process the input and add points to a canonical 3D space by incrementally updating a scene’s latent state. Spann3R [62] introduces a spatial memory network, enabling multiview input and improving efficiency to eliminate the need for global alignment. MUSt3R [63] extends the DUSt3R architecture by introducing a symmetric design and a memory mechanism, effectively reducing computational complexity when handling multiview inputs. Closely, CUT3R [3] proposes a Continuous Updating Transformer that simultaneously updates the state with new information and retrieves the information stored in the state. This formulation is general and able to handle both video and photo collections, and process both static and dynamic scenes. However, with the increased number of processed frames, memory-based methods face capacity constraints, which can result in the degradation or loss of information from earlier frames. To address this issue, Point3R [64] takes inspiration from the human memory mechanism and proposes a spatial pointer memory, where each pointer is anchored at a 3D position and links to a dynamically evolving spatial feature. Besides, Driv3R [65] extends the memory mechanism to support efficient temporal integration, enabling large-scale dynamic scene reconstruction from sequences of multiview inputs.
@@
In addition, several methods are proposed to develop new SfM pipelines for efficient 3D reconstruction. Specifically, Light3R-SfM [66] replaces optimization-based global alignment with a learnable latent alignment module, enabling the efficient SfM and 3D reconstruction. Regist3R [67] introduces a stereo foundation model to build a scalable incremental SfM pipeline for efficient 3D reconstruction.
@@
To facilitate accurate 3D reconstruction, Pow3R [68] flexibly integrates available priors at test time, such as camera intrinsics, sparse or dense depth, or relative poses, as lightweight and diverse conditioning. In contrast, Rig3R [69] exploits the rig metadata as conditions to improve both the camera pose estimation and 3D reconstruction. In addition, MoGe [70] replaces the scale-invariant pointmaps used in DUSt3R with the affine-invariant pointmaps, enabling superior geometry learning. It additionally introduces a novel global alignment solver to improve the geometry accuracy by addressing the scale and shift issues in the reconstructed affine-invariant pointmaps. Test3R [71] takes advantage of test time training to improve the geometric consistency of pointmaps. AerialMegaDepth [72] instead focuses on aerial-ground geometric reconstruction from a data perspective.
@@
3D Gaussian Splatting (GS) [18] is a recent advance for efficient 3D reconstruction and rendering built on rasterization. 3DGS is a point-based representation that each point is associated with geometry attributes (i.e. center position, shape, orientation, and opacity $\alpha$) and Spherical Harmonics (SH) appearance attributes. Despite its high fidelity in reconstruction, 3DGS requires per-scene optimization, which limits its training efficiency and generalization capabilities. Recently, feed-forward 3DGS reconstruction methods have been developed, leveraging neural networks to directly predict Gaussian parameters. These approaches eliminate the need for per-scene optimization and enable generalizable novel view synthesis. We categorize these methods based on the representation of predicted Gaussian outputs: image, volume, triplane, and pointmap.
@@
Figure 5: Representative frameworks with different outputs of 3D Gaussian representations, including Gaussian image, Gaussian volume, Gaussian triplane, and Gaussian PointMap. The samples are adapted from [74] [75] [76] [77].
2.3.1 Gaussian Image
A Gaussian image refers to a 2D image-based representation of 3D Gaussians, where each pixel encodes a 3D Gaussian. As a pioneering effort, Splatter Image [74] employs a U-Net encoder-decoder architecture [78] to predict pixel-aligned 3D Gaussians for single-view 3D object reconstruction as illustrated in Fig. 5(a).

### v3

@@ Abstract
Figure 1: Timeline of representative feed-forward works categorized by methods.
Figure 2: The above figure includes various applications of feed-forward models in 3D reconstruction and Novel View Synthesis (NVS). The samples are adapted from [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
@@ 1 Introduction
3D reconstruction and rendering are long-standing and central challenges in computer vision and computer graphics. They enable a wide range of applications, from digital content creation, augmented reality, and virtual reality to robotics, autonomous systems, and digital twins. Traditionally, high-quality 3D reconstruction and view synthesis have relied on optimization-based pipelines such as Structure-from-Motion (SfM) [13] and 3D Gaussian Splatting [14]. However, these methods typically require per-scene optimization, converge slowly, and rely on dense multiview inputs, limiting their practicality in real-world scenarios. In light of this, feed-forward methods have emerged as an important research line in 3D vision.
@@
Feed-forward models are learning-based methods that infer outputs, such as 3D geometry or novel views, in a single forward pass of a neural network, without requiring per-scene iterative optimization in classical methods. Early examples include cost-volume-based Multi-View Stereo [15] and layered representations such as multiplane images (MPI) [16], which demonstrate the potential of learning-based inference over per-scene optimization. In recent years, fueled by breakthroughs in deep learning and neural representations, feed-forward methods [17, 18, 19] have emerged as a transformative alternative in 3D reconstruction and view synthesis. By learning from large scale data, these models enable orders of magnitude faster inference and better generalization compared to classical optimization methods. Their ability to produce predictions in real time makes them particularly attractive for time-sensitive and scalable applications, such as robotic perception and interactive 3D asset creation.
@@
This survey focuses on feed-forward methods developed primarily after the emergence of neural radiance fields (NeRF) [20] in 2020, which catalyzed a rapid evolution in feed-forward approaches, as shown in Fig. 1. We present a comprehensive review of feed-forward methods for 3D reconstruction and view synthesis, with an emphasis on the scene representations, core architectures, and downstream applications that define this fast-evolving area. Specifically, we systematically categorize existing approaches based on their underlying scene representations, which determine how 3D structure and appearance are modeled and rendered. In this way, we identify five major categories: 1) models built on Neural Radiance Fields (NeRF) [20], which leverage volumetric rendering through learned radiance fields; 2) pointmap-based approaches [1], which operate on pixel-aligned 3D pointmaps; 3) 3D Gaussian Splatting (3DGS)-based models [14], which use rasterizable Gaussian primitives for fast and efficient rendering; 4) methods based on other 3D representations like mesh, occupancy and signed distance function (SDF); and 5) 3D-free models, which leverage deep neural networks to synthesize views directly without a explicit 3D representation. For each category, we provide an in-depth analysis of representative and state-of-the-art methods, highlighting their core architectural designs and feature representations.
@@
We also highlight several high-impact 3D vision applications enabled by feed-forward methods, as illustrated in Fig. 2. These methods offer scalable, fast, and generalizable solutions across domains, supporting tasks such as multi-view 3D reconstruction, dynamic 3D Gaussian Splatting, and diffusion-based view synthesis. Beyond reconstruction and view synthesis, they also advance image matching, 3D-aware segmentation, and optical flow estimation. In the broader context of robotics and SLAM, feed-forward models enable efficient real-time scene understanding and tracking, while in digital humans, they support generalizable avatar reconstruction from sparse inputs.
@@
Despite impressive progress, feed-forward models still face major challenges, including limited modality diversity in datasets, poor generalization in free-viewpoint synthesis, and the high computational cost of long-context processing. Addressing these challenges will require advances in efficient architectures and scalable datasets. Finally, we conclude with the societal impact of this method, highlighting the importance of responsible deployment and transparent modeling practices.
@@ 2 Methods
We broadly categorize the feed-forward 3D reconstruction and view synthesis methods into five categories based on their underlying representation: NeRF models (Sec. 2.1), Pointmap models (Sec. 2.2), 3DGS models (Sec. 2.3), models employing other common representations (e.g., mesh, occupancy, SDFs in Sec. 2.4), and 3D-free models (Sec. 2.5).
Figure 3: Representative frameworks of feed-forward NeRF, including methods based on (a) 1D features, (b) 2D features, and (c) & (d) 3D features. The samples are adapted from [21] [22] [23] [10].
@@ 2.1 NeRF
Neural radiance fields (NeRF) [20] have recently gained significant attention for high-quality novel view synthesis using volumetric scene representations and differentiable volume rendering. By leveraging MLPs, NeRF reconstructs 3D scenes from multiview 2D images, enabling the generation of novel views with excellent multiview consistency. However, a major limitation of NeRF is its requirement for per-scene optimization, so it cannot generalize to unseen scenes. To address this, feed-forward approaches have been proposed, where neural networks learn to infer NeRF representations directly from sparse input views, thereby eliminating the need for scene-specific optimization. As a pioneering feed-forward NeRF work, PixelNeRF [19] introduces a conditional NeRF framework that leverages pixel-aligned image features extracted from input images, allowing the model to generalize across diverse scenes and perform novel view synthesis from sparse observations. A large number of follow-ups adopt various techniques for feed-forward NeRF, and we broadly categorize them into the following categories based on feature representations.
@@
Several methods have been proposed to encode a global 1D latent code for NeRF prediction, where the same latent code is shared between all 3D points in a scene. For example, CodeNeRF [21], as illustrated in Fig. 3(a), introduces a disentanglement strategy that jointly learns separate embeddings for texture and shape, along with an MLP conditioned on these embeddings to predict the color and volumetric density of each 3D point. ShaRF [24] introduces latent codes for shape and appearance, which serve as conditioning inputs for NeRF reconstruction. In addition, Shap-E [25] encodes point clouds and RGBA input images into a series of latent vectors that are subsequently used for generating NeRF’s MLPs.
@@
2D feature-based methods typically leverage an image encoder to extract image features of source views and obtain features of arbitrary 3D points by ray projection without relying on 3D intermediate features. For example, GRF [22], as illustrated in Fig. 3(b) projects each 3D point along a camera ray onto source views to extract corresponding multiview features. These features are then aggregated and passed through an MLP to predict color and density. IBRNet [26] follows a similar approach, projecting 3D points onto nearby source views to extract image features that are aggregated across views for NeRF inference. NeRFormer [27] also employs ray-projected features and performs multiview feature aggregation to guide the NeRF prediction. In addition, SRF [28] projects 3D points onto multiview images to construct a stereo feature matrix, which is processed by a 2D CNN to produce view-aligned features for radiance prediction. To provide additional geometric cues for NeRF prediction, GNT [29] introduces a view transformer that uses epipolar constraint to aggregate projected features from multiple views in a geometrically consistent manner. Its successor, GNT-MOVE [30], bridges the view transformer with the Mixture-of-Experts concept from large language models, enhancing its cross-scene generalization capability. Instead of relying on epipolar lines, MatchNeRF [31] explicitly models the correspondence information by computing the similarity between ray-projected features from pairs of nearby source views, using this information as a conditioning input for NeRF prediction. In addition, ContraNeRF [32] introduces geometry-aware contrastive learning [33] to learn multiview consistent 2D features to mitigate synthetic-to-real generalization issues.
@@
3D Volume Features. Inspired by multiview stereo (MVS) [15, 34, 35], MVSNeRF [23] constructs 3D cost volumes that store the matching costs of pixels/features across multiple images from input images (usually along depth values). These cost volumes are used to generate a neural scene encoding volume, which stores per-voxel features representing both local geometry and appearance, as shown in Fig. 3(c). For any 3D point, its features are obtained via trilinear interpolation from the encoding volume and then decoded by an MLP to predict the corresponding density and color. To improve rendering quality in both fine-detail areas and occluded regions, GeoNeRF [36] extends MVSNeRF by first constructing cascaded 3D cost volumes for each source view, followed by an attention-based volume aggregation across views. Similarly addressing occlusions, NeuRay [37] leverages cost volumes to predict the visibility of 3D points, which can identify feature inconsistencies caused by occlusion. WaveNeRF [38] further enhances geometry reconstruction by integrating wavelet frequency volumes into the MVS pipeline to preserve high-frequency details.
For efficient rendering, ENeRF [39] samples a limited number of points near the scene surface, guided by coarse scene geometry predicted from a cascaded cost volume, thereby speeding up rendering; Instead of constructing cost volumes for all reference input views, MuRF [40] efficiently constructs a target view frustum volume to aggregate information from the input images; To improve the quality of geometry estimation, GeFu [41] introduces an adaptive cost aggregation module that learns to reweight contributions from different source views, enabling more accurate cost volume construction.
@@
3D Triplane Features. The 3D triplane serves as an efficient volumetric representation [42, 43], making it highly compatible with feed-forward models. Specifically, Large Reconstruction Model (LRM) [10] employs a large transformer-based encoder-decoder architecture and directly regresses a feature triplane representation as shown in Fig. 3(d), enabling NeRF prediction from triplane features. Pf-LRM [44] extends LRM to a pose-free setting, which jointly reconstructs the triplane NeRF representations and predicts relative camera poses. TripoSR [45] further advances LRM by carefully curating a subset of Objaverse [46] for training, along with improved architectures and training strategies. To address the scarcity, licensing constraints, and inherent biases of 3D data, LRM-Zero [47] introduces the Zeroverse dataset and performs training entirely on synthesized data. Beyond reconstruction alone, several methods integrate LRM with diffusion models. For example, Instant3D [48] first leverages a fine-tuned 2D diffusion model [49] to generate 4-view images from a text prompt and then uses a transformer-based large reconstruction model to predict a NeRF; DMV3D [50] incorporates LRM into multiview diffusion, which gradually reconstructs a clean triplane NeRF representation from noisy multiview images in the diffusion process.
@@
In addition to the methods discussed above, several approaches have explored feed-forward NeRF reconstruction with other types of features. For example, VisionNeRF [51] proposes to leverage vision transformer [52] and convolutional networks to extract global 1D features and 2D image features, respectively, and constructs a multi-level feature map that serves as the conditioning inputs of NeRF prediction to enhance rendering quality, particularly in occluded regions. MINE [53] integrates NeRF and multiplane image (MPI) [54] representations to enable generalizable, occlusion-aware 3D reconstruction from a single image.
@@ 2.2 Pointmap
Figure 4: The framework of feed-forward pointmap reconstruction. It also supports broader tasks, such as camera estimation, depth estimation, and point tracking, within the same framework. The samples are adapted from [3].
@@
To handle more views, Fast3R [59] builds on DUSt3R and designs a global fusion transformer to process multiview inputs simultaneously. MV-DUSt3R [60] instead leverages multiview decoder blocks to learn both the reference-to-source and source-to-source view relationships, thereby extending DUSt3R to a multiview setting. SLAM3R [61] introduces an Image-to-Points module that enables simultaneous processing of multiview inputs, effectively enhancing reconstruction quality without sequential reconstruction.
@@
Notably, several works incorporate a memory mechanism that incrementally processes inputs and updates a scene’s latent state, progressively adding points to a canonical 3D representation. Spann3R [62] introduces a spatial memory network that enables multiview inputs while improving efficiency, removing the need for global alignment. Similarly, MUSt3R [63] extends the DUSt3R architecture with a symmetric design and a memory mechanism, effectively reducing computational complexity when processing multiview inputs. CUT3R [3] proposes a Continuous Updating Transformer that simultaneously updates the state with new information and retrieves the information stored in the state. This formulation is general, supporting video and photo collections as well as static and dynamic scenes. However, with the increased number of processed frames, memory-based methods face capacity constraints, which can result in the degradation or loss of information from earlier frames. To address this issue, Point3R [64] takes inspiration from the human memory mechanism and proposes a spatial pointer memory, where each pointer is anchored at a 3D position and links to a dynamically evolving spatial feature. Driv3R [65] extends the memory mechanism to support efficient temporal integration, enabling large-scale dynamic scene reconstruction from multiview input sequences.
@@
yuelei here In addition, several methods are proposed to develop new SfM pipelines for efficient 3D reconstruction. Specifically, Light3R-SfM [66] replaces optimization-based global alignment with a learnable latent alignment module, enabling the efficient SfM and 3D reconstruction. Regist3R [67] introduces a stereo foundation model to build a scalable incremental SfM pipeline for efficient 3D reconstruction.
@@
To facilitate accurate 3D reconstruction, Pow3R [68] flexibly integrates available priors at test time, such as camera intrinsics, sparse or dense depth, or relative poses, as lightweight and diverse conditioning. In contrast, Rig3R [69] exploits the rig metadata as conditions to improve both the camera pose estimation and 3D reconstruction. Meanwhile, MoGe [70] improves geometry learning by replacing DUSt3R’s scale-invariant pointmaps with affine-invariant ones, and further introduces a global alignment solver to correct scale and shift errors in affine-invariant pointmaps. Similarly focused on geometric consistency, Test3R [71] takes advantage of test time training to improve the geometric consistency of pointmaps. AerialMegaDepth [72] instead focuses on aerial-ground geometric reconstruction from a data perspective.
@@
3D Gaussian Splatting (GS) [14] is a recent advance for efficient 3D reconstruction and rendering built on rasterization. Despite its high fidelity in reconstruction, 3DGS requires per-scene optimization, which limits its training efficiency and generalization capabilities. Recently, feed-forward 3DGS reconstruction methods have been developed by leveraging neural networks to directly predict Gaussian parameters, eliminating the need for per-scene optimization. We categorize these methods based on the representation of predicted Gaussian outputs: 2D map and 3D volume.
@@
Figure 5: Representative frameworks with different outputs of 3D Gaussian representations, including Gaussian Map (e.g., Image Splatter and pointmap) and Gaussian Volume (e.g., Triplane and 3D Volume). The samples are adapted from [74] [75] [76] [77].
2.3.1 Gaussian Map